### v2

# Relatives of the Hermitian curve

Abstract
We introduce the notion of a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z)=0$
with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties together with various examples. One of the basic properties is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
Key Words: Plane curve, Finite field, Rational point
MSC: 14G15, 14H50, 14G05, 11G20
1 Our object
Throughout this paper, $q$ denotes an even power of a prime number, so $\sqrt{q}$ is also a power of the prime number.
A Hermitian curve is a plane curve defined by
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}=0,$ (1)
where $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition ${}^{t}\!A^{(\sqrt{q})}=A$. Here the symbol $A^{(\sqrt{q})}$ means taking $\sqrt{q}$-th power on each entry of $A$.
Forgetting the Hermitian condition, the equation (1) still defines a nonsingular plane curves of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which will be denoted by $C_{A}$. We call such a curve a relative of the Hermitian curve, or shortly, a Hermitian-relative curve, and want to study them.
All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [7, Proof of Proposition 3.7]. We will give a proof of this fact in the next section for the convenience of readers.
Any Hermitian-relative curve $C$ over $\mathbb{F}_{q}$ has a particular property on the number $N_{q}(C)$ of $\mathbb{F}_{q}$-points:
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
The proof of this fact is also given in the following section.
Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$. By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and the image of the element in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably. So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
Remark 1.1
If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition ${}^{t}\overline{A}^{(\sqrt{q})}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
Proof.
Let $A$ be any source of $\overline{A}$. Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that ${}^{t}A^{(\sqrt{q})}=\lambda A$. Taking $\sqrt{q}$-th power and transpose of the both side, $A=\lambda^{\sqrt{q}}\left({}^{t}\!A^{(\sqrt{q})}\right)$. So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$. Therefore there is an element $\rho\in\mathbb{F}_{q}$ so that $\rho^{1-\sqrt{q}}=\lambda$ (which is a special case of Hilbert’s Theorem 90). Then $\overline{\rho A}=\overline{A}$ and $\rho A$ satisfies the Hermitian condition. ∎
Annotation
Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, hereafter, for any $\mathbb{F}_{q}$-point, we understand all values of the coordinates are elements of $\mathbb{F}_{q}$.
2 Basic properties
The polynomial of type (1) is a linear combination of the monomials
$x^{\sqrt{q}+1},y^{\sqrt{q}+1},z^{\sqrt{q}+1},x^{\sqrt{q}}y,y^{\sqrt{q}}z,z^{\sqrt{q}}x,xy^{\sqrt{q}},yz^{\sqrt{q}},zx^{\sqrt{q}}$
over $\mathbb{F}_{q}$, and vice versa. So those polynomials forms an $\mathbb{F}_{q}$-vector space of dimension $9$. Moreover the vector space is stable under the linear transformations on coordinates $x,y,z$. We should note that by a linear transformation
$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
with $T\in PGL(3,\mathbb{F}_{q})$, the polynomial $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ goes to
$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$.
For given two matrices $A,B\in GL(3,\mathbb{F}_{q})$ (or in $PGL(3,\mathbb{F}_{q})$), we say that $A$ and $B$ are equivalent if two curves $C_{A}$ and $C_{B}$ are projectively equivalent over $\mathbb{F}_{q}$. This equivalence relation is denoted by $A\sim B$. In other words, $A\sim B$ if there is a matrix $T\in GL(3,\mathbb{F}_{q})$ and $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $\lambda B={}^{t}T^{(\sqrt{q})}AT$ (or $B={}^{t}T^{(\sqrt{q})}AT$ in $PGL(3,\mathbb{F}_{q})$).
Remark 2.1
Even if $\det A=0$, the curve defined by (1) makes sense; however the curve is nonsingular if and only if $\det A\neq 0$. Actually, the partial derivatives of $f=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ are given by
$(f_{x},f_{y},f_{z})=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A.$ (2)
Proposition 2.2
For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is projectively equivalent to $C_{A^{(\sqrt{q})}}$ over $\mathbb{F}_{q}$.
Proof.
Let $u,v,w$ be the dual coordinate of the dual projective plane with point-line incidence $ux+vy+wz=0$. By (2), the equation after eliminating $x,y,z$ from
$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (3)
and (1) is an equation of the dual curve. Taking entry-wise $\sqrt{q}$-th power of (1),
$(x^{q},y^{q},z^{q})A^{(\sqrt{q})}\begin{pmatrix}x^{\sqrt{q}}\\ y^{\sqrt{q}}\\ z^{\sqrt{q}}\end{pmatrix}=0,$
which can be written
$(u^{\sqrt{q}},v^{\sqrt{q}},w^{\sqrt{q}}){}^{t}\!A^{-1}\begin{pmatrix}u\\ v\\ w\end{pmatrix}=0$
by (3). On the other hand, since
(${}^{t}\!\left({}^{t}\!A\right)^{(\sqrt{q})}\left({}^{t}\!A^{-1}\right){}^{t}\!A=A^{(\sqrt{q})},$
${}^{t}\!A^{-1}\sim A^{(\sqrt{q})}$. This completes the proof. ∎
Next is the first main purpose of this section.
Theorem 2.3
Let $f(x,y,z)\in\mathbb{F}_{q}[x,y,z]$ be a homogeneous polynomial of degree $\sqrt{q}+1$. The curve $C$ defined by $f(x,y,z)=0$ in $\mathbb{P}^{2}$ is projectively equivalent over $\overline{\mathbb{F}}_{q}$ to a Hermitian curve of degree $\sqrt{q}+1$ if and only if $C$ is a Hermitian-relative curve, that is, there exists $A\in PGL(3,\mathbb{F}_{q})$ such that
$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Proof.
(Pardini) First suppose that $C$ is projectively equivalent to
$g(x,y,z)=x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$
over $\overline{\mathbb{F}}_{q}$. Then there is a matrix $T=(t_{i,j})\in GL(3,\overline{\mathbb{F}}_{q})$ so that
$f(x,y,z)=g((x,y,z){}^{t}T),$
that is,
$f(x,y,z)$ $=\sum_{j=1}^{3}\left(t_{j1}x+t_{j2}y+t_{j3}z\right)^{\sqrt{q}+1}$
$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
Note that each entry of ${}^{t}T^{(\sqrt{q})}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So ${}^{t}T^{(\sqrt{q})}T\in PGL(3,\mathbb{F}_{q})$.
To see the converse, let us consider a morphism associated with an assigned $A\in PGL(3,\mathbb{F}_{q})$:
$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto{}^{t}T^{(\sqrt{q})}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
If $\Phi_{A}(S)=\Phi_{A}(T)$ for $S,T\in PGL(3,\overline{\mathbb{F}}_{q})$, then
(${}^{t}(ST^{-1})^{(\sqrt{q})}A(ST^{-1})=A$
which means $ST^{-1}$ gives a linear automorphism of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$.
The automorphism group of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ is finite if $\sqrt{q}>2$, because the genus of $C_{A}$ is grater than $1$. When $\sqrt{q}=2$, $C_{A}$ is an elliptic curve over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$. The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite. Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [6, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$. Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that ${}^{t}T^{(\sqrt{q})}AT=B$. In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve. ∎
In the next corollary, $T_{P}(C)$ denotes the (embedded) tangent line to a plane curve $C$ at $P$, and $i(C.T_{P}(C);P)$ the intersection multiplicity of $C$ and $T_{P}(C)$ at $P$.
Corollary 2.4
Suppose that $q>4$. Among nonsingular plane curves of degree $\sqrt{q}+1$, a Hermitian-relative curve $C$ is characterized by the property: for any $\overline{\mathbb{F}}_{q}$-point $P$ of $C$, $i(C.T_{P}(C);P)$ is either $\sqrt{q}$ or $\sqrt{q}+1$.
Proof.
This characterization comes from [4, Theorem 6.1, and Theorem 3.5] together with Theorem 2.3
The following is the second main purpose of this section.
Theorem 2.5
Let $C$ be a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$. Then
$N_{q}(C)\equiv 1\mod\sqrt{q}.$
In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points on $C$ is not empty.
Proof.
Let $g=\frac{1}{2}\sqrt{q}(\sqrt{q}-1)$, which is the genus of any Hermitian-relative curve. Let $C^{\prime}$ be a Hermitian curve of degree $\sqrt{q}+1$, e.g., $C^{\prime}$ is one defined by $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0.$ By Theorem 2.3, $C$ and $C^{\prime}$ are projectively equivalent over some extension field $\mathbb{F}_{q^{s}}$. Let
$Z_{C}(t)=\prod_{j=1}^{2g}(1-\alpha_{j}t)/(1-t)(1-qt)$
be the zeta function of $C$ over $\mathbb{F}_{q}$. By Weil’s theorem,
$N_{q^{s}}(C)=1+q^{s}-\sum_{j=1}^{2g}\alpha_{j}^{s}\text{\ and \ }\alpha_{j}=\sqrt{q}\eta_{j}\text{\ with \ }|\eta_{j}|=1.$ (4)
On the other hand, since $C^{\prime}$ is a maximal curve over $\mathbb{F}_{q}$, its zeta function is
$Z_{C^{\prime}}(t)=\left(1-(-\sqrt{q})t\right)^{2g}/(1-t)(1-qt).$
Since $C$ and $C^{\prime}$ are isomorphic over $\mathbb{F}_{q^{s}}$,
$\{\alpha_{1}^{s},\dots,\alpha_{2g}^{s}\}=\{\overbrace{(-\sqrt{q})^{s},\dots,(-\sqrt{q})^{s}}^{2g}\}.$
(As for this classical theory, one may consult [9, Chapter 5].)
Since we may choose $s$ to be even, $\eta_{j}^{s}=1$ for each $j=1,\dots,2g$. In particular, $\eta_{j}$ is an algebraic integer, and so is $\sum_{j=1}^{2g}\eta_{j}$. Going back to the formula (4) (with $s=1$),
$\sum_{j=1}^{2g}\eta_{j}=\frac{q+1-N_{q}(C)}{\sqrt{q}}.$
Since $\sqrt{q}$ is an integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
$N_{q}(C)=q+1-\sqrt{q}(\sum_{j=1}^{2g}\eta_{j})\equiv 1\mod\sqrt{q}.$
This completes the proof. ∎
Corollary 2.6
For a relative of the Hermitian curve $C$ over $\mathbb{F}_{q}$, there is a nonnegative integer $m=m_{C}$ with
$m=q\text{\ or \ }\sqrt{q}+2\geq m\geq 0$
so that $N_{q}(C)=m\sqrt{q}+1$. Moreover $m_{C}=q$ if and only if either $C$ is Hermitian, or $q=4$ and $C$ is projectively equivalent to the curve
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{2}z^{\sqrt{q}+1}=0.$
over $\mathbb{F}_{4}$, where $\omega\in\mathbb{F}_{4}\setminus\mathbb{F}_{2}$.
Proof.
The additional statement is somewhat classical [3, 8, 5]. For the first statement, when $q=4$, there is nothing new from Theorem 2.5. Hence we assume that $q>4$. Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$. Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A\begin{pmatrix}x_{0}\\ y_{0}\\ z_{0}\end{pmatrix}=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{{}^{t}\!A^{(\sqrt{q})}}$. If $C$ is not Hermitian, $C_{A}$ and $C_{{}^{t}\!A^{(\sqrt{q})}}$ are distinct irreducible curves. Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$. ∎
3 Example I - diagonal case
Remaining three sections, selected examples are explained; however, the division of sections is rather superficial (cf. Example 3.4 and Example 4.1). In order to compute examples, some notes will be explained. The trace $\mathbb{F}_{q}\ni x\mapsto x^{\sqrt{q}}+x\in\mathbb{F}_{\sqrt{q}}$ and the norm $\mathbb{F}_{q}^{\ast}\ni x\mapsto x^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are indicated by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
Lemma 3.1
There are two exact sequences:
$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}}-\alpha&&&\end{array};$ (5)
$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}-1}&&&\end{array},$ (6)
where $i$ means natural inclusion.
Proof.
This lemma is a special case of Hilbert’s Theorem 90, and the proof is easy. ∎
Corollary 3.2
  1. (i)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}$
    $X^{\sqrt{q}}-X-\beta=0$ (7)
[…]
  2. (ii)
    The equation of $X$ with $\beta\in\mathbb{F}_{q}^{\ast}$
    $X^{\sqrt{q}-1}-\beta=0$ (8)
[…]
Proof.
(i) Let $\alpha\in\mathbb{F}_{q}$ be a root of (7). Then $\beta=\alpha^{\sqrt{q}}-\alpha$. Hence $\operatorname{Tr}\beta=0$ by (5). Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (5) again. When $\alpha$ is a root of (7), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
(ii) can be proved by using (6). ∎
In this section, we handle only diagonal matrices; $\operatorname{diag}[a,b,c]$ denotes the diagonal matrix whose diagonal entries are $a,b,c$ from the upper left to the lower right.
Let $A\in PGL(3,\mathbb{F}_{q})$, which is diagonal. Then $A$ is equivalent to one of the following three types.
Example 3.3
(Type 0) $A\sim\operatorname{diag}[1,1,1]$. So the corresponding curve $C_{A}$ is projectively equivalent to the Hermitian curve $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$, and $N_{q}(C_{A})=(\sqrt{q})^{3}+1$. Hence $m_{C_{A}}=q$.
Example 3.4
(Type 1) $A\sim\operatorname{diag}[1,1,\omega]$ for some $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. Hence $C_{A}$ is projectively equivalent to the curve $D_{\omega}$ defined by
$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\omega z^{\sqrt{q}+1}=0.$ (9)
If $(x_{0},y_{0},z_{0})\in(\mathbb{F}_{q})^{3}$ is a solution of (9), then it is a solution of the simultaneous equations:
[…]
because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}$ and $\{1,\omega\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
$D_{\omega}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}.$
Hence $N_{q}(C_{A})=\sqrt{q}+1$, and $m_{C_{A}}=1$.
Example 3.5
(Type 2) $A\sim\operatorname{diag}[1,\omega,\omega^{\prime}]$ for some $\omega,\omega^{\prime}\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$ with $\omega(\omega^{\prime})^{-1}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$. Hence $C_{A}$ is projectively equivalent the curve $D_{\omega,\omega^{\prime}}$ defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{\prime}z^{\sqrt{q}+1}=0.$ (10)
Since any two elements of $\{1,\omega,\omega^{\prime}\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, there is a unique relation up to $\mathbb{F}_{\sqrt{q}}^{\ast}$-multiple among those three elements, say
$\alpha+\beta\omega+\gamma\omega^{\prime}=0.$
Since none of three coefficients $\alpha,\beta,\gamma$ are 00, we may assume that $\alpha=1$, and the first coordinate of any solution of (10) is also $1$. Then $D_{\omega,\omega^{\prime}}(\mathbb{F}_{q})=\{(1,\zeta,\eta)|\operatorname{Nm}\zeta=\beta,\operatorname{Nm}\eta=\gamma\}$. Hence $N_{q}(C_{A})=(\sqrt{q}+1)^{2}$, and hence $m_{C}=\sqrt{q}+2$.
4 Example II - non-diagonal case
The following examples, including the diagonal case explained above, are inspired by Hirschfeld’s classification of cubic plane curves over any finite fields [2, Section 11]. Actually, the family of supersingular plane cubic curves over $\mathbb{F}_{4}$ coincides with the family of relatives of the Hermitian curve over $\mathbb{F}_{4}$, because any curves in each family is characterized by disappearance of the term $xyz$ in its equation (see, [1, IV, Prop. 4.21]).
Example 4.1
Let $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, and
$A=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}.$
Then $C_{A}$ is defined by
$x^{\sqrt{q}}y+xy^{\sqrt{q}}+\omega z^{\sqrt{q}+1}=0.$
The curve $C_{A}$ is projectively equivalent over $\mathbb{F}_{q}$ to a diagonal curve of type 1. In fact, choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$. Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then it is easy to see that
${}^{t}\!T_{0}^{(\sqrt{q})}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\omega].$
Example 4.2
Let us consider the matrix $A=\begin{pmatrix}1&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. So the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (11)
Proposition 4.3
For the curve (11), $C_{A}(\mathbb{F}_{q})=\{(0,0,1)\}$, and hence $m_{C_{A}}=0$.
Proof.
For any $(x_{0},y_{0},z_{0})\in C_{A}(\mathbb{F}_{q})$, three elements $x_{0}^{\sqrt{q}+1}$, $y_{0}^{\sqrt{q}+1}$, and $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})$ are in $\mathbb{F}_{\sqrt{q}}$. Since $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, we have
$x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=0\textrm{\ and \ }y_{0}^{\sqrt{q}+1}=0.$
Hence $x_{0}=y_{0}=0$. ∎
Example 4.4
Next we consider the matrix $A=\begin{pmatrix}\omega&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The corresponding curve $C_{A}$ is defined by
$\omega x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (12)
Proposition 4.5
For the curve (12),
$\begin{split}C_{A}(\mathbb{F}_{q})=\{P_{\zeta}=(\zeta,1,0)|&\zeta^{\sqrt{q}+1}=-1\}\cup\{R=(0,0,1)\}\\ &\cup\{Q_{\xi\rho}=(\xi,\rho,1)|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.\end{split}$
In particular, $N_{q}(C_{A})=q+\sqrt{q}+1$ and $m_{C_{A}}=\sqrt{q}+1$.
Proof.
On the line $\{z=0\}$ [^1] During computation of examples, frequently use this type of abbreviation., the equation is $\omega(x^{\sqrt{q}+1}+y^{\sqrt{q}+1})=0.$ Hence
$C_{A}(\mathbf{F}_{q})\cap\{z=0\}=\{P_{\zeta}|\zeta^{\sqrt{q}+1}=-1.\}$
In affine space $\mathbb{P}^{2}\setminus\{z=0\}$, we may assume that the $z$-coordinate is always $1$. If $(x_{0},y_{0},1)$ is an $\mathbb{F}_{q}$-point of $C_{A}$, it satisfies
$\omega(x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1})+y_{0}^{\sqrt{q}}+y_{0}=0,$ (13)
and vice versa. Since $x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1}$ and $y_{0}^{\sqrt{q}}+y_{0}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, (13) is equivalent to
[…]
Hence
$C_{A}(\mathbb{F}_{q})\cap\{z\neq 0\}=\{R\}\cup\{Q_{\xi\rho}|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.$
It is easy to show that |$|\{P_{\xi}\}|=\sqrt{q}+1$, and |$|\{R\}\cup\{Q_{\xi\rho}\}|=1+(\sqrt{q}-1)(\sqrt{q}+1).$
Example 4.6
Let $A=\begin{pmatrix}1&0&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ Hence the curve $C_{A}$ is defined by
$x^{\sqrt{q}+1}+xy^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$
Proposition 4.7
For the curve $C_{A}$ in Example 4.6, we have
$N_{q}(C_{A})=q+1,\text{\ and so\ }m_{C_{A}}=\sqrt{q}.$
Proof.
First consider the case $x=0$. Then the equation becomes $y^{\sqrt{q}}z+yz^{\sqrt{q}}=0$, which has obviously $\sqrt{q}+1$ homogeneous solutions that are $\mathbb{F}_{q}$points.
Next suppose that $x=1$. Then the equation becomes
$01+y^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=$ 0 . (14)
Let $(y_{0},z_{0})$ be any $\mathbb{F}_{q}$-solution of (14). Since $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})\in\mathbb{F}_{\sqrt{q}}$, the remaining term $1+y_{0}^{\sqrt{q}}$ is also an element of $\mathbb{F}_{\sqrt{q}}$, and so is $y_{0}^{\sqrt{q}}$. Hence $y_{0}=y_{0}^{q}=(y_{0}^{\sqrt{q}})^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}}$. Therefore the point $(y_{0},z_{0})$ satisfies
$01+y_{0}+(z_{0}+z_{0}^{\sqrt{q}})y_{0}=$ 0 (15)
from (14). Put $z_{0}+z_{0}^{\sqrt{q}}=t_{0}$, which is an elements of $\mathbb{F}_{\sqrt{q}}$. If $t_{0}\neq-1$, $y_{0}=\frac{-1}{t_{0}+1}$, which is uniquely determined by $t_{0}$. For each $t_{0}\in\mathbb{F}_{\sqrt{q}}\setminus\{-1\}$, we can find $\sqrt{q}$ solutions of $z+z^{\sqrt{q}}=t_{0}$ in $\mathbb{F}_{q}$. If $t_{0}=-1$, There are no solutions of (15) on $y_{0}$. Hence $|C_{A}(\mathbb{F}_{q})\cap\{x\neq 0\}|=\sqrt{q}(\sqrt{q}-1).$ Therefore the number of $\mathbb{F}_{q}$-points of $C_{A}$ is totally
$(\sqrt{q}+1)+\sqrt{q}(\sqrt{q}-1)=q+1.$
This completes the proof. ∎
5 Example III - non-diagonal case (2)
In this section, we handle the curve
$x^{\sqrt{q}}z+xy^{\sqrt{q}}+\omega yz^{\sqrt{q}}=0,$ (16)
where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. The number of $\mathbb{F}_{q}$-points of curves in each examples explained in previous two sections does not depend on the choice of $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, however that of the curve (16) depends on the choice of $\omega$. So we denote the corresponding matrix $\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&\omega&0\end{pmatrix}$ by $A_{\omega}$.
Lemma 5.1
Let $C_{A_{\omega}}$ be a curve defined by (16).
(i)   $C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z=0\}=\{(1,0,0),(0,1,0)\}.$
${\rm(ii)}\begin{split}(C_{A_{\omega}}(\mathbb{F}_{q})\cap&\{z\neq 0\})\cap\{y^{\sqrt{q}+1}\neq 1\}\\ &=\left\{\left(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1},y_{0},1\right)\mid y_{0}\in\mathbb{F}_{q},\ y_{0}^{\sqrt{q}+1}-1\neq 0\right\}.\end{split}$
In particular,
$N_{q}(C_{A_{\omega}})\geq q-\sqrt{q}+1.$
Proof.
The assertion (i) is obvious. Let $(x_{0},y_{0},1)\in C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\}$. Then
$x_{0}^{\sqrt{q}}+x_{0}y_{0}^{\sqrt{q}}+\omega y_{0}=0.$ (17)
Taking $\sqrt{q}$-th power of this equation, we have
$x_{0}+x_{0}^{\sqrt{q}}y_{0}+\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}=0.$ (18)
[…]
$x_{0}(y_{0}^{\sqrt{q}+1}-1)-(\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2})=0.$
Hence, when $y_{0}^{\sqrt{q}+1}-1\neq 0$, $x_{0}=\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1}$. Conversely, by straightforward computation, we get
$(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})^{\sqrt{q}}+(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})y_{0}^{\sqrt{q}}+\omega y_{0}=0.$
Putting (i), (ii) together, we have $q-(\sqrt{q}+1)+2$ points over $\mathbb{F}_{q}$. ∎
For computing the number of $\mathbb{F}_{q}$-points of $(C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\})\cap\{y^{\sqrt{q}+1}=0\},$ we fix a primitive element $\omega_{0}\in\mathbb{F}_{q}^{\ast}$.
Note that $\omega_{0}^{s}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ if and only if $s\equiv 0\mod\sqrt{q}+1$.
Lemma 5.2
(1) Suppose that $q$ is odd. Then
(i) $-1=\omega_{0}^{\frac{q-1}{2}},$ and
(ii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{\frac{\sqrt{q}+1}{2}+s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}.$
(2) Suppose that $q$ is even. Then
(iii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}$.
Proof.
(i) Since $\omega_{0}$ is of order $q-1$ in the multiplicative group $\mathbb{F}_{q}^{\ast}$ and $q-1$ is even, $\omega_{0}^{\frac{q-1}{2}}=-1$.
(ii) For $b\in\mathbb{F}_{q}^{\ast}$, $\operatorname{Tr}b=0$ means $b^{\sqrt{q}-1}=-1$. Letting $b=\omega_{0}^{l}$, we have $\omega_{0}^{l(\sqrt{q}-1)}=\omega_{0}^{\frac{q-1}{2}}$. Hence $l(\sqrt{q}-1)\equiv\frac{\sqrt{q}+1}{2}(\sqrt{q}-1)\mod q-1$, which means that $l\equiv\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1$.
(iii) Similar to the case (ii). ∎
Theorem 5.3
  1. (1)
    If $(\sqrt{q}+1,3)=1$, then
    $N_{q}(C_{A_{\omega}})=q+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}.$
  2. (2)
    When $\sqrt{q}+1$ is divisible by $3$,
    1. (a)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q+2\sqrt{q}+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}+2,\text{\ and}$
    2. (b)
      if $\omega^{\frac{\sqrt{q}+1}{3}}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
      $N_{q}(C_{A_{\omega}})=q-\sqrt{q}+1\text{\ and \ }m_{C_{A_{\omega}}}=\sqrt{q}-1.$
Proof.
We have to count the number of $\mathbb{F}_{q}$-solutions of the simultaneous equations
[…]
By changing the sign $y\mapsto-y,$ this system is equivalent to
[…]
For a given $y_{0}\in\mathbb{F}_{q}^{\ast}$ with $y_{0}^{\sqrt{q}+1}=1$, we ask whether the equation on $x$
$x^{\sqrt{q}}-y_{0}^{\sqrt{q}}x-\omega y_{0}=0$ (19)
has an $\mathbb{F}_{q}$-solution or not. Since $\operatorname{Nm}y_{0}^{\sqrt{q}}=\operatorname{Nm}y_{0}=1$, there is an element $c\in\mathbb{F}_{q}^{\ast}$ such that $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$ by Lemma 3.1. Therefore the problem above is transformed whether the equation on $cx$
$(cx)^{\sqrt{q}}-(cx)-\omega c^{\sqrt{q}}y_{0}=0$ (20)
has an $\mathbb{F}_{q}$-solution or not. From Corollary 3.2, (20) has an $\mathbb{F}_{q}$-solution if and only if $\operatorname{Tr}(\omega c^{\sqrt{q}}y_{0})=0$, and in this case all solutions of (20) are $\mathbb{F}_{q}$-solutions.
Let $i,j,k$ be integers with $11\leq i,j,k\leq q-$ 1 satisfying
$\omega=\omega_{0}^{k},\ c=\omega_{0}^{j},\text{\ and \ }y_{0}=\omega_{0}^{i}.$
Since $y_{0}^{\sqrt{q}+1}=1$, there is a unique integer $l$ with $11\leq l\leq\sqrt{q}+$ 1 so that $i\equiv l(\sqrt{q}-1)\mod q-1$. Since $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$,
$j(1-\sqrt{q})\equiv l(\sqrt{q}-1)\sqrt{q}\mod q-1.$
Hence
$j\equiv-l\sqrt{q}\mod\sqrt{q}+1.$ (21)
Since $\omega c^{\sqrt{q}}y_{0}=\omega_{0}^{k+j\sqrt{q}+l(\sqrt{q}-1)},$ $\operatorname{Tr}\omega c^{\sqrt{q}}y_{0}=0$ if and only if
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}\end{cases}$ (22)
by Lemma 5.2. On the other hand,
$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv k-l(q-\sqrt{q}+1)\mod\sqrt{q}+1$ (23)
by (21). Together with $\sqrt{q}\equiv-1\mod\sqrt{q}+1$ and $q\equiv-\sqrt{q}\equiv 1\mod\sqrt{q}+1,$ (22) and (23) imply
$k-3l\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}.\end{cases}$ (24)
Therefore the number of $y_{0}$’s for each of which (19) has $\mathbb{F}_{q}$-solutions, in fact $\sqrt{q}$-solutions, coincides with the number of $l$’s with $11\leq l\leq\sqrt{q}+$ 1 each of which satisfies (24) for the fixed $k$.
Let us consider the map
$\begin{array}[]{rccc}\varphi:&\mathbb{Z}/(\sqrt{q}+1)&\to&\mathbb{Z}/(\sqrt{q}+1)\\ &x&\mapsto&3x\end{array}.$
If $\sqrt{q}+1$ is not divisible by $3$, $\varphi$ is bijective. Hence there exists a unique $l$ satisfying (24) for any fixed $k$. If $3$ is a factor of $\sqrt{q}+1$, then
$|\varphi^{-1}(m)|=\begin{cases}3&\text{if $m\in\operatorname{Im}\varphi$}\\ 0&\text{else}.\end{cases}$
Hence if $k\equiv 0\mod 3$, there are three $l$’s satisfying (24), and otherwise no $l$’s. The condition $k\equiv 0\mod 3$ is equivalent to the condition $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}.$ This completes the proof. ∎
Finally we pose a question: when $q\geq 16$, does a Hermitian-relative curve $C$ with $22\leq m_{C}\leq\sqrt{q}-$ 2 exist?
Acknowledgment. The second author was partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (2022R1A2C1012291).

### v3

@@
 $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z)=0$
-with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties together with various examples.
-One of the basic properties is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
+with $A\in GL(3,\mathbb{F}_{q})$, and study their basic properties, one of which is that the number of $\mathbb{F}_{q}$-points of any relative of the Hermitian curve of degree $\sqrt{q}+1$ is congruent to $1$ modulo $\sqrt{q}$.
+In the latter part of this paper, we classify those curves having two or more rational inflexions.
 Key Words: Plane curve, Finite field, Rational point
 MSC: 14G15, 14H50, 14G05, 11G20
-1 Our object
-Throughout this paper, $q$ denotes an even power of a prime number, so $\sqrt{q}$ is also a power of the prime number.
-A Hermitian curve is a plane curve defined by
-$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}=0,$ (1)
-where $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition ${}^{t}\!A^{(\sqrt{q})}=A$. Here the symbol $A^{(\sqrt{q})}$ means taking $\sqrt{q}$-th power on each entry of $A$.
+1 Introduction
+Throughout this paper, $q$ denotes an even power of a prime number $p$.
+So $\sqrt{q}$ is also a power of $p$.
+For $A\in GL(3,\mathbb{F}_{q})$, $A^{\ast}$ denotes the matrix ${}^{t}\!A^{(\sqrt{q})}$, where the superscript $(\sqrt{q})$ indicates taking entry-wise $\sqrt{q}$-th power, and the superscripted prefix $t$ taking transpose.
+The Hermitian curve is a plane curve defined by
+$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ (1)
+with the condition $A=A^{\ast}$
+In this case, we say that $A$ satisfies the Hermitian condition.
 Forgetting the Hermitian condition, the equation (1) still defines a nonsingular plane curves of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$, which will be denoted by $C_{A}$.
 We call such a curve a relative of the Hermitian curve, or shortly, a Hermitian-relative curve, and want to study them.
-All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [7, Proof of Proposition 3.7].
+All Hermitian-relative curves, including Hermitian curves themselves, are projectively equivalent each other over the algebraic closure $\overline{\mathbb{F}}_{q}$ of $\mathbb{F}_{q}$, which is due to Pardini [5, Proof of Proposition 3.7].
 We will give a proof of this fact in the next section for the convenience of readers.
@@
 Since the equation (1) is homogeneous, the curve $C_{A}$ depends only on the image $\overline{A}$ of $A$ under the canonical map $GL(3,\mathbb{F}_{q})\to PGL(3,\mathbb{F}_{q})$.
-By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and the image of the element in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably.
+By abuse of notation, $\overline{A}$ will be frequently denoted by $A$, namely an element of $GL(3,\mathbb{F}_{q})$ and its image in $PGL(3,\mathbb{F}_{q})$ will be used interchangeably.
 So the equation $A=B$ in $PGL(3,\mathbb{F}_{q})$ means that there exists an element $\lambda\in\mathbb{F}_{q}^{\ast}$ so that $A=\lambda B$ in $GL(3,\mathbb{F}_{q})$.
-Remark 1.1
-If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition ${}^{t}\overline{A}^{(\sqrt{q})}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
+Remark 1.1.
+If $\overline{A}\in PGL(3,\mathbb{F}_{q})$ satisfies the Hermitian condition, that is to say, $\overline{A^{\ast}}=\overline{A}$, then we can choose an element $A\in GL(3,\mathbb{F}_{q})$ satisfying the Hermitian condition whose image in $PGL(3,\mathbb{F}_{q})$ is the assigned $\overline{A}$.
 Proof.
 Let $A$ be any source of $\overline{A}$.
-Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that ${}^{t}A^{(\sqrt{q})}=\lambda A$.
-Taking $\sqrt{q}$-th power and transpose of the both side, $A=\lambda^{\sqrt{q}}\left({}^{t}\!A^{(\sqrt{q})}\right)$.
+Then there is an element $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $A^{\ast}=\lambda A$.
+Taking $\ast$-operation on the both side, $A=\lambda^{\sqrt{q}}A^{\ast}$.
 So $A=\lambda^{\sqrt{q}+1}A$, which implies $\lambda^{\sqrt{q}+1}=1$.
@@
 ∎
-Annotation
-Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, hereafter, for any $\mathbb{F}_{q}$-point, we understand all values of the coordinates are elements of $\mathbb{F}_{q}$.
-2 Basic properties
 The polynomial of type (1) is a linear combination of the monomials
@@
 We should note that by a linear transformation
-$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
+$\begin{pmatrix}x\\ y\\ z\end{pmatrix}\mapsto T^{-1}\begin{pmatrix}x\\ y\\ z\end{pmatrix}$
 with $T\in PGL(3,\mathbb{F}_{q})$, the polynomial $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}$ goes to
-$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
-The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$.
-For given two matrices $A,B\in GL(3,\mathbb{F}_{q})$ (or in $PGL(3,\mathbb{F}_{q})$), we say that $A$ and $B$ are equivalent if two curves $C_{A}$ and $C_{B}$ are projectively equivalent over $\mathbb{F}_{q}$.
-This equivalence relation is denoted by $A\sim B$.
-In other words, $A\sim B$ if there is a matrix $T\in GL(3,\mathbb{F}_{q})$ and $\lambda\in\mathbb{F}_{q}^{\ast}$ such that $\lambda B={}^{t}T^{(\sqrt{q})}AT$ (or $B={}^{t}T^{(\sqrt{q})}AT$ in $PGL(3,\mathbb{F}_{q})$).
-Remark 2.1
+$(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}AT\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
+In other words, two relatives of the Hermitian curve $C_{A}$ and $C_{B}$, are projectively equivalent over $\mathbb{F}_{q}$ if and only if there is a certain matrix $T\in PGL(3,\mathbb{F}_{q})$ such that $B=T^{\ast}AT$ in $PGL(3,\mathbb{F}_{q})$.
+This is sometimes denoted by $A\sim B$.
+Remark 1.2.
 Even if $\det A=0$, the curve defined by (1) makes sense; however the curve is nonsingular if and only if $\det A\neq 0$.
@@
 $(f_{x},f_{y},f_{z})=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A.$ (2)
-Proposition 2.2
-For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is projectively equivalent to $C_{A^{(\sqrt{q})}}$ over $\mathbb{F}_{q}$.
-Proof.
-Let $u,v,w$ be the dual coordinate of the dual projective plane with point-line incidence $ux+vy+wz=0$. By (2), the equation after eliminating $x,y,z$ from
-$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (3)
-and (1) is an equation of the dual curve. Taking entry-wise $\sqrt{q}$-th power of (1),
-$(x^{q},y^{q},z^{q})A^{(\sqrt{q})}\begin{pmatrix}x^{\sqrt{q}}\\ y^{\sqrt{q}}\\ z^{\sqrt{q}}\end{pmatrix}=0,$
-which can be written
-$(u^{\sqrt{q}},v^{\sqrt{q}},w^{\sqrt{q}}){}^{t}\!A^{-1}\begin{pmatrix}u\\ v\\ w\end{pmatrix}=0$
-by (3). On the other hand, since
-(${}^{t}\!\left({}^{t}\!A\right)^{(\sqrt{q})}\left({}^{t}\!A^{-1}\right){}^{t}\!A=A^{(\sqrt{q})},$
-${}^{t}\!A^{-1}\sim A^{(\sqrt{q})}$. This completes the proof. ∎
-Next is the first main purpose of this section.
-Theorem 2.3
+The properties mentioned above hold true after any field extension $\mathbb{F}_{q}$.
+As is explained in the next section, the tangent line at a point of $C_{A}$ contacts of multiplicity $\sqrt{q}$ in general, but at some special points of multiplicity $\sqrt{q}+1$.
+In the letter case, the point is called an inflexion.
+In Section 3, we give the classification up to $PGL(3,\mathbb{F}_{q})$ of the curves $C_{A}$ with two or more inflexions defined over $\mathbb{F}_{q}$.
+In the classification, three types appear and it can be summarized in Table 1.
+Table 1: Classification
+$N_{q}(C_{A})$ inflexions classes
+(a) $\sqrt{q}^{3}+1$ $\sqrt{q}^{3}+1$ 1
+(b) $\sqrt{q}+1$ $\sqrt{q}+1$ $\sqrt{q}$
+(c) $q+1$ $2$ $\frac{1}{2}(\sqrt{q}+1)(\sqrt{q}-2)$
+In the table, the first column indicates the label of each type which agrees with the item in Theorem 3.1, the second the number of $\mathbb{F}_{q}$-points of $C_{A}$ belongs the assigned type, the third the number of inflexions, and the fourth is the number of $\mathbb{F}_{q}$-equivalent classes in each type.
+Annotation.
+$\bullet$ Although a homogeneous coordinates of a point has an ambiguity by nonzero multiple, we understand all values of the coordinates of any $\mathbb{F}_{q}$-point are elements in $\mathbb{F}_{q}$.
+Moreover, we use two words “$\mathbb{F}_{q}$-point” and “rational point” interchangeably.
+$\bullet$ The set of $\mathbb{F}_{q}$-points of a curve $C$ is denoted by $C(\mathbb{F}_{q})$.
+$\bullet$ For a field, the multiplicative group of nonzero elements is marked by the superscript $\ast$.
+$\bullet$ The trace map from $\mathbb{F}_{q}$ to $\mathbb{F}_{\sqrt{q}}$, that is, $\mathbb{F}_{q}\ni\alpha\mapsto\alpha+\alpha^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}},$ and the norm map $\mathbb{F}_{q}^{\ast}$ to $\mathbb{F}_{\sqrt{q}}^{\ast}$, that is, $\mathbb{F}_{q}^{\ast}\ni\alpha\mapsto\alpha^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are simply denoted by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
+$\bullet$ Unless we specify the field of definition, ‘projectively equivalent’ means ‘projectively equivalent over $\mathbb{F}_{q}$’.
+$\bullet$ Sometimes a curve defined by $g=0$ will be indicated by $\{g=0\}$.
+2 Basic properties
+We shall start this section with Pardini’s observation [5], which is our motivation to study the Hermitian-relative curve.
+Theorem 2.1.
 Let $f(x,y,z)\in\mathbb{F}_{q}[x,y,z]$ be a homogeneous polynomial of degree $\sqrt{q}+1$.
 The curve $C$ defined by $f(x,y,z)=0$ in $\mathbb{P}^{2}$ is projectively equivalent over $\overline{\mathbb{F}}_{q}$ to a Hermitian curve of degree $\sqrt{q}+1$ if and only if $C$ is a Hermitian-relative curve, that is, there exists $A\in PGL(3,\mathbb{F}_{q})$ such that
-$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
+$f(x,y,z)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A{}^{t}\!(x,y,z).$
 Proof.
@@
 $f(x,y,z)$ $=\sum_{j=1}^{3}\left(t_{j1}x+t_{j2}y+t_{j3}z\right)^{\sqrt{q}+1}$
-$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}}){}^{t}T^{(\sqrt{q})}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
-Note that each entry of ${}^{t}T^{(\sqrt{q})}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So ${}^{t}T^{(\sqrt{q})}T\in PGL(3,\mathbb{F}_{q})$.
+$=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})T^{\ast}T\begin{pmatrix}x\\ y\\ z\end{pmatrix}.$
+Note that each entry of $T^{\ast}T$ gives a coefficient of distinct monomial of $f(x,y,z)$. So $T^{\ast}T\in PGL(3,\mathbb{F}_{q})$.
 To see the converse, let us consider a morphism associated with an assigned $A\in PGL(3,\mathbb{F}_{q})$:
-$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto{}^{t}T^{(\sqrt{q})}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
+$\Phi_{A}:PGL(3,\overline{\mathbb{F}}_{q})\ni T\mapsto T^{\ast}AT\in PGL(3,\overline{\mathbb{F}}_{q}).$
 If $\Phi_{A}(S)=\Phi_{A}(T)$ for $S,T\in PGL(3,\overline{\mathbb{F}}_{q})$, then
-(${}^{t}(ST^{-1})^{(\sqrt{q})}A(ST^{-1})=A$
+$(ST^{-1})^{\ast}A(ST^{-1})=A$
 which means $ST^{-1}$ gives a linear automorphism of $C_{A}$ over $\overline{\mathbb{F}}_{q}$ in $\mathbb{P}^{2}$.
@@
 The linear automorphism group of any elliptic curve in $\mathbb{P}^{2}$ is also finite.
-Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [6, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$.
-Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that ${}^{t}T^{(\sqrt{q})}AT=B$.
+Therefore the image of $\Phi_{A}$ is a constructible set (which is a theorem of Chevalley [4, Corollary 2 in I §8]) of dimension $8$ in the target space $PGL(3,\overline{\mathbb{F}}_{q})$.
+Therefore two images of $\Phi_{A}$ and $\Phi_{B}$ meet in $PGL(3,\overline{\mathbb{F}}_{q})$ for any two $A,B\in PGL(3,\mathbb{F}_{q})$, that is, there is an element $T\in PGL(3,\overline{\mathbb{F}}_{q})$ such that $T^{\ast}AT=B$.
 In particular, if one takes as $B=I_{3}$, one gets a projective transformation over $\overline{\mathbb{F}}_{q}$ which sends $C_{A}$ to a Hermitian curve.
@@
 In the next corollary, $T_{P}(C)$ denotes the (embedded) tangent line to a plane curve $C$ at $P$, and $i(C.T_{P}(C);P)$ the intersection multiplicity of $C$ and $T_{P}(C)$ at $P$.
-Corollary 2.4
+Corollary 2.2.
 Suppose that $q>4$.
@@
 Proof.
-This characterization comes from [4, Theorem 6.1, and Theorem 3.5] together with Theorem 2.3
-The following is the second main purpose of this section.
-Theorem 2.5
+This characterization comes from [2, Theorem 6.1, and Theorem 3.5] together with Theorem 2.1
+The following is the main purpose of this section.
+Theorem 2.3.
 Let $C$ be a relative of the Hermitian curve of degree $\sqrt{q}+1$ over $\mathbb{F}_{q}$. Then
 $N_{q}(C)\equiv 1\mod\sqrt{q}.$
-In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points on $C$ is not empty.
+In particular, the set $C(\mathbb{F}_{q})$ of $\mathbb{F}_{q}$-points is not empty.
 Proof.
@@
 $\{\alpha_{1}^{s},\dots,\alpha_{2g}^{s}\}=\{\overbrace{(-\sqrt{q})^{s},\dots,(-\sqrt{q})^{s}}^{2g}\}.$
-(As for this classical theory, one may consult [9, Chapter 5].)
+(As for this classical theory, one may consult [7, Chapter 5].)
 Since we may choose $s$ to be even, $\eta_{j}^{s}=1$ for each $j=1,\dots,2g$.
@@
 $\sum_{j=1}^{2g}\eta_{j}=\frac{q+1-N_{q}(C)}{\sqrt{q}}.$
-Since $\sqrt{q}$ is an integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
+Since $\sqrt{q}$ is a rational integer by our assumption, $\sum_{j=1}^{2g}\eta_{j}$ is a rational number. Therefore it is a rational integer. Hence
 $N_{q}(C)=q+1-\sqrt{q}(\sum_{j=1}^{2g}\eta_{j})\equiv 1\mod\sqrt{q}.$
 This completes the proof. ∎
-Corollary 2.6
+Corollary 2.4.
 For a relative of the Hermitian curve $C$ over $\mathbb{F}_{q}$, there is a nonnegative integer $m=m_{C}$ with
@@
 Let $C=C_{A}$, and $(x_{0},y_{0},z_{0})\in C(\mathbb{F}_{q})$.
-Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A\begin{pmatrix}x_{0}\\ y_{0}\\ z_{0}\end{pmatrix}=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{{}^{t}\!A^{(\sqrt{q})}}$.
-If $C$ is not Hermitian, $C_{A}$ and $C_{{}^{t}\!A^{(\sqrt{q})}}$ are distinct irreducible curves.
+Then $(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{0},y_{0},z_{0})=0.$ Taking $\sqrt{q}$-th power, and noting $x_{0}^{q}=x_{0},y_{0}^{q}=y_{0},z_{0}^{q}=z_{0}$, we know that $(x_{0},y_{0},z_{0})\in C_{A^{\ast}}$.
+If $C$ is not Hermitian, $C_{A}$ and $C_{A^{\ast}}$ are distinct irreducible curves.
 Hence the intersection of these two curves consists of at most $(\sqrt{q}+1)^{2}$ points, which gives an upper bound for $N_{q}(C)$.
 ∎
-3 Example I - diagonal case
-Remaining three sections, selected examples are explained; however, the division of sections is rather superficial (cf.
-Example 3.4 and Example 4.1).
-In order to compute examples, some notes will be explained.
-The trace $\mathbb{F}_{q}\ni x\mapsto x^{\sqrt{q}}+x\in\mathbb{F}_{\sqrt{q}}$ and the norm $\mathbb{F}_{q}^{\ast}\ni x\mapsto x^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ are indicated by $\operatorname{Tr}$ and $\operatorname{Nm}$ respectively.
-Lemma 3.1
+Before closing this section, we shall digress a little to explain a couple of geometric properties of relatives of the Hermitian curve.
+Proposition 2.5.
+For $A\in GL(3,\mathbb{F}_{q})$, the dual curve of $C_{A}$ is $C_{{}^{t}\!A^{-1}}$.
+Proof.
+Let $u,v,w$ be the dual coordinate of the dual projective plane with point-line incidence $ux+vy+wz=0$. By (2), the equation after eliminating $x,y,z$ from
+$(u,v,w)=(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A$ (4)
+and (1) is an equation of the dual curve. Taking entry-wise $\sqrt{q}$-th power of (1),
+$(x^{q},y^{q},z^{q})A^{(\sqrt{q})}\begin{pmatrix}x^{\sqrt{q}}\\ y^{\sqrt{q}}\\ z^{\sqrt{q}}\end{pmatrix}=0,$
+which can be written
+$(u^{\sqrt{q}},v^{\sqrt{q}},w^{\sqrt{q}}){}^{t}\!A^{-1}\begin{pmatrix}u\\ v\\ w\end{pmatrix}=0$
+by (4). This completes the proof. ∎
+For a Hermitian-relative curve $C_{A}$, the curve $C_{A^{\ast}}$ is called the mirror curve of $C_{A}$.
+Of course, the mirror curve of $C_{A^{\ast}}$ is the original one $C_{A}$, and $C_{A}$ is Hermitian if and only if $C_{A^{\ast}}=C_{A}$.
+Proposition 2.6.
+For a mirror pair $(C_{A},C_{A^{\ast}})$, the following properties hold.
+  1. (1)
+    $C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$.
+  2. (2)
+    For $P\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$, $P$ is an inflexion of $C_{A}$ if and only if an inflexion of $C_{A^{\ast}}$. In this case, $T_{P}(C_{A})=T_{P}(C_{A}^{\ast})$.
+  3. (3)
+    For a non-inflexion $P\in C_{A}(\mathbb{F}_{q})$, then there is a point $P^{\prime}\in C_{A}(\mathbb{F}_{q})$ so that $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$.
+In this case, $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P$.
+In particular $P^{\prime}$ is a non-inflexion of $C_{A^{\ast}}$.
+Proof.
+If $C_{A}$ is Hermitian, then the assertions (1) and (2) are trivially true, and (3) is nonsense. So $C_{A}$ may be assumed to be non-Hermitian.
+We already saw the assertion (1) in the proof of Corollary 2.4.
+(2) Since $(T^{\ast}AT)^{\ast}=T^{\ast}A^{\ast}T$ for any $T\in PGL(3,\mathbb{F}_{q})$, we may assume that $P=(0,0,1)$ and $T_{P}C_{A}=\{y=0\}$.
+Let $A=(a_{ij})$.
+Since $P\in C_{A}$, $a_{33}=0$.
+Hence an affine equation of $C_{A}$ on $\{z\neq 0\}$ is given by
+$a_{11}x^{\sqrt{q}+1}+(a_{12}y+a_{13})x^{\sqrt{q}}+(a_{21}x+a_{22}y+a_{23})y^{\sqrt{q}}+a_{31}x+a_{32}y=0.$
+Since $T_{P}C_{A}=\{y=0\}$ and the origin $P$ is an inflexion of $C_{A}$, we have $a_{31}=0$, $a_{13}=0$ and $a_{11}\neq 0$. Hence we may assume that $a_{11}=1$. To sum up,
+$A=\begin{pmatrix}1&a_{12}&0\\ a_{21}&a_{22}&a_{23}\\ 0&a_{32}&0\end{pmatrix}.$
+Hence
+$A^{\ast}=\begin{pmatrix}1&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&a_{32}^{\sqrt{q}}\\ 0&a_{23}^{\sqrt{q}}&0\end{pmatrix},$
+in other words, an affine equation of $C_{A^{\ast}}$ is
+$x^{\sqrt{q}+1}+a_{21}^{\sqrt{q}}yx^{\sqrt{q}}+(a_{12}^{\sqrt{q}}x+a_{22}^{\sqrt{q}}y+a_{32}^{\sqrt{q}})y^{\sqrt{q}}+a_{32}^{\sqrt{q}}y=0.$
+This shows that the origin is also an inflexion of $C_{A^{\ast}}$ with tangent line $\{y=0\}$.
+(3) By Corollary 2.2, $C_{A}.T_{P}C_{A}=\sqrt{q}P+P^{\prime}$ for some point $P^{\prime}$.
+Since two divisors $C_{A}.T_{P}C_{A}$ and $\sqrt{q}P$ are defined over $\mathbb{F}_{q}$, so is the residual one $P^{\prime}$.
+Hence $P^{\prime}\in C_{A}(\mathbb{F}_{q})=C_{A^{\ast}}(\mathbb{F}_{q})$.
+Let $P=(x_{0},y_{0},z_{0})$ and $P^{\prime}=(x_{1},y_{1},z_{1})$.
+Then
+$(x_{0}^{\sqrt{q}},y_{0}^{\sqrt{q}},z_{0}^{\sqrt{q}})A{}^{t}\!(x_{1},y_{1},z_{1})$
+by Remark 2. Hence
+$(x_{1}^{\sqrt{q}},y_{1}^{\sqrt{q}},z_{1}^{\sqrt{q}})A^{\ast}{}^{t}\!(x_{0},y_{0},z_{0}),$
+which means $P\in C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}$, and hence $C_{A^{\ast}}.T_{P^{\prime}}C_{A^{\ast}}=\sqrt{q}P^{\prime}+P.$
+3 Hermitian-relative curves with two or more rational inflexions
+The purpose of this section is to give the classification of the Hermitian-relative curves having two or more rational inflexions, which is summarized in the following theorem.
+Theorem 3.1.
+Let $C$ be a relative of the Hermitian curve. Suppose that $C$ has at least two rational inflexions. Then $C$ is projectively equivalent to
+$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$ (5)
+[…]
+  1. (a)
+[…]
+  2. (b)
+[…]
+    $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (6)
+[…]
+    For two curves $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ and $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta^{\prime}z^{\sqrt{q}+1}=0$ with $\eta,\,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ are projectively equivalent to each other if and only if $\eta\eta^{\prime-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
+Therefore, there are exactly $\sqrt{q}$ non-equivalent Hermitian-relative curves, each of which has $\sqrt{q}+1$ rational points and all of them are inflexions.
+  3. (c)
+[…]
+3.1 Some equations
+Since any trace map and norm map between finite fields are surjective, it is easy to show Hilbert’s Theorem 90 in our case:
+Lemma 3.2.
 There are two exact sequences:
-$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}}-\alpha&&&\end{array};$ (5)
-$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{\sqrt{q}-1}&&&\end{array},$ (6)
+$\begin{array}[]{rcccccccl}0\to&\mathbb{F}_{\sqrt{q}}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}&\to&\mathbb{F}_{q}&\xrightarrow{\operatorname{Tr}}&\mathbb{F}_{\sqrt{q}}&\to 0\\ &&&\alpha&\mapsto&\alpha-\alpha^{\sqrt{q}}&&&\end{array}$ (7)
+$\begin{array}[]{rcccccccl}1\to&\mathbb{F}_{\sqrt{q}}^{\ast}&\stackrel{{\scriptstyle i}}{{\to}}&\mathbb{F}_{q}^{\ast}&\to&\mathbb{F}_{q}^{\ast}&\xrightarrow{\operatorname{Nm}}&\mathbb{F}_{\sqrt{q}}^{\ast}&\to 1\\ &&&\alpha&\mapsto&\alpha^{1-\sqrt{q}}&&&\end{array},$ (8)
 where $i$ means natural inclusion.
-Proof.
-This lemma is a special case of Hilbert’s Theorem 90, and the proof is easy. ∎
-Corollary 3.2
+Corollary 3.3.
   1. (i)
@@
 Then $\beta=\alpha^{\sqrt{q}}-\alpha$.
-Hence $\operatorname{Tr}\beta=0$ by (5).
-Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (5) again.
+Hence $\operatorname{Tr}\beta=0$.
+Conversely, if $\operatorname{Tr}\beta=0$, there is an element $\alpha\in\mathbb{F}_{q}$ such that $\alpha^{\sqrt{q}}-\alpha=\beta$ by (7).
 When $\alpha$ is a root of (9), the set of all roots is $\{\alpha+\gamma|\gamma\in\mathbb{F}_{\sqrt{q}}\}$.
 (ii) can be proved by using (8). ∎
-In this section, we handle only diagonal matrices; $\operatorname{diag}[a,b,c]$ denotes the diagonal matrix whose diagonal entries are $a,b,c$ from the upper left to the lower right.
-Let $A\in PGL(3,\mathbb{F}_{q})$, which is diagonal. Then $A$ is equivalent to one of the following three types.
-Example 3.3
-(Type 0) $A\sim\operatorname{diag}[1,1,1]$.
-So the corresponding curve $C_{A}$ is projectively equivalent to the Hermitian curve $x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+z^{\sqrt{q}+1}=0$, and $N_{q}(C_{A})=(\sqrt{q})^{3}+1$.
-Hence $m_{C_{A}}=q$.
-Example 3.4
-(Type 1) $A\sim\operatorname{diag}[1,1,\omega]$ for some $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. Hence $C_{A}$ is projectively equivalent to the curve $D_{\omega}$ defined by
-$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\omega z^{\sqrt{q}+1}=0.$ (9)
+The above corollary can be generalized slightly as follows:
+Proposition 3.4.
+Let $\alpha\in\mathbb{F}_{q}^{\ast}$ and $\beta\in\mathbb{F}_{q}$, and consider the equation in $X$:
+$X^{\sqrt{q}}+\alpha X+\beta=0.$ (11)
+  1. (i)
+[…]
+  2. (ii)
+[…]
+Proof.
+For a fixed $\alpha\in\mathbb{F}_{q}^{\ast}$, consider the $\mathbb{F}_{\sqrt{q}}$-linear map $\varphi_{\alpha}:\mathbb{F}_{q}\to\mathbb{F}_{q}$ given by $\varphi_{\alpha}(z)=z^{\sqrt{q}}+\alpha z$.
+Since $\operatorname{Nm}(-\alpha)=\operatorname{Nm}\alpha$, the kernel of $\varphi_{\alpha}$ is nontrivial if and only if $\operatorname{Nm}\alpha=1$ by Corollary 3.3 (ii).
+Therefore if $\operatorname{Nm}\alpha\neq 1$, then $\varphi_{\alpha}$ is an isomorphism.
+Hence there is a unique element of $\mathbb{F}_{q}$, say $z_{0}$, so that $\varphi_{\alpha}(z_{0})=-\beta$.
+So the proof of (ii) is done.
+Next suppose that $\operatorname{Nm}\alpha=1$.
+In this case, $\dim\operatorname{Ker}\varphi_{\alpha}>0$ from the first paragraph of this proof, in particular there is an element $z_{0}\in\mathbb{F}_{q}^{\ast}$ such that $z_{0}^{\sqrt{q}-1}+\alpha=0$.
+Hence
+$\operatorname{Ker}\varphi_{\alpha}=\{\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{0\}.$
+If (11) has a solution $z_{1}$, the set of all solutions of (11) is
+$\{z_{1}+\zeta z_{0}\mid\zeta\in\mathbb{F}_{\sqrt{q}}^{\ast}\}\cup\{z_{1}\},$
+which is a subset of $\mathbb{F}_{q}$. So the additional statement of (i) is done.
+Now we prove the first part of (i).
+To prove the ‘only if’ part, we may suppose that $\beta\neq 0$.
+Let $z_{1}$ be an $\mathbb{F}_{q}$-solution of (11).
+Then $\alpha=-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}}$.
+Since $\operatorname{Nm}\alpha=1$ and $(z_{1}^{\sqrt{q}})^{\sqrt{q}}=z_{1}$, we have
+11 11 $=$ $\operatorname{Nm}\alpha=-(\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})^{\sqrt{q}}(-\frac{z_{1}^{\sqrt{q}}+\beta}{z_{1}})$
+$=$ $\frac{z_{1}^{\sqrt{q}+1}+\beta^{\sqrt{q}}z_{1}^{\sqrt{q}}+\beta z_{1}+\beta^{\sqrt{q}+1}}{z_{1}^{\sqrt{q}+1}}.$
+Hence
+$z_{1}^{\sqrt{q}}+\beta^{1-\sqrt{q}}z_{1}+\beta=0.$
+Since $z_{1}$ satisfies (11), we get $\alpha=\beta^{1-\sqrt{q}}$.
+Conversely, if $\beta=0$, obviously (11) has an $\mathbb{F}_{q}$-solution.
+Hence we may consider the case $\alpha=\beta^{1-\sqrt{q}}$ with $\beta\neq 0$.
+In this case, (11) is
+$X^{\sqrt{q}}+\beta^{1-\sqrt{q}}X+\beta=0$ (12)
+which is equivalent to
+$(\beta X)^{\sqrt{q}}+(\beta X)+\beta^{\sqrt{q}+1}=0.$ (13)
+Since
+$(-\beta^{\sqrt{q}+1})^{\sqrt{q}-1}=\beta^{q-1}=1$
+because $\beta\in\mathbb{F}_{q}^{\ast}$, we know $-\beta^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
+Since $\operatorname{Tr}:\mathbb{F}_{q}\to\mathbb{F}_{\sqrt{q}}$ is surjective, the equation (13) in $(\beta X)$ has a solution in $\mathbb{F}_{q}$, and also (12) has.
+∎
+3.2 Proof of Theorem 3.1
+Lemma 3.5.
+Let $C$ be a relative of the Hermitian curve with at least two rational inflexions. Then we can choose a system of coordinates $x,y,z$ of $\mathbb{P}^{2}$ over $\mathbb{F}_{q}$ such that
+  • $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions of $C$, and
+  • $T_{P}(C)=\{y=0\}$ and $T_{Q}(C)=\{x=0\}$.
+Proof.
+Since $C.T_{P}(C)=(\sqrt{q}+1)P$ and $C.T_{Q}(C)=(\sqrt{q}+1)Q$, three $\mathbb{F}_{q}$-lines $T_{P}(C)$, $T_{Q}(C)$ and $PQ$ are not concurrent, where $PQ$ denotes the line through $P$ and $Q$.
+Therefore, we may choose coordinates as $T_{P}(C)=\{y=0\}$, $T_{Q}(C)=\{x=0\}$ and $PQ=\{z=0\}$.
+∎
+Lemma 3.6.
+Under the situation described in Lemma 3.5, the curve is defined by $(x^{\sqrt{q}},y^{\sqrt{q}},z^{\sqrt{q}})A\,{}^{t}\!(x,y,z)=0,$ where
+$A=\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&a_{33}\end{pmatrix}\in PGL(3,\mathbb{F}_{q})\ \text{with}\ a_{12}a_{21}a_{33}\neq 0.$
+In other words, $C$ is defined by
+$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$ (14)
+Proof.
+Let $C=C_{A}$, where $A=(a_{ij})\in PGL(3,\mathbb{F}_{q})$. Since $(1,0,0)A{}^{t}\!(1,0,0)=0$ and $(0,1,0)A{}^{t}\!(0,1,0)=0$, $a_{11}=a_{22}=0$.
+Note that the tangent line at $(b_{1},b_{2},b_{3})\in C$ is given by
+$(b_{1}^{\sqrt{q}},b_{2}^{\sqrt{q}},b_{3}^{\sqrt{q}})A{}^{t}\!(x,y,z)=0.$
+Hence $a_{13}=0$ because $T_{P}(C)=\{y=0\}$, and $a_{23}=0$ because $T_{Q}(C)=\{x=0\}$. Therefore an equation of $C$ is of the form
+$a_{12}x^{\sqrt{q}}y+a_{21}xy^{\sqrt{q}}+a_{31}xz^{\sqrt{q}}+a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0.$
+Since $P$ is an inflexion of $C$ and $T_{P}(C)=\{y=0\}$, we have $C.\{y=0\}=(\sqrt{q}+1)P$. Hence the equation
+$a_{31}xz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0$
+has the root at $z=0$ of multiplicity $(\sqrt{q}+1)$.
+So $a_{31}=0$.
+Similarly, since $C.\{x=0\}=(\sqrt{q}+1)Q$, from the equation $a_{32}yz^{\sqrt{q}}+a_{33}z^{\sqrt{q}+1}=0,$ we know $a_{32}=0$.
+Finally $\det A\neq 0$ implies $a_{12}a_{21}a_{33}\neq 0$.
+∎
+Corollary 3.7.
+After a suitable projective transformation over $\mathbb{F}_{q}$, $C$ is defined by
+$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0$
+with $\omega\in\mathbb{F}_{q}^{\ast}$. Conversely, a curve defined by this type of equation is Hermitian-relative with two or more inflexions.
+Proof.
+Since $a_{33}\neq 0$, we may suppose that $a_{33}=1$ in (14). If one consider $a_{12}y$ as new $y$, then the equation becomes
+$x^{\sqrt{q}}y+a_{21}a_{12}^{-{\sqrt{q}}}xy^{\sqrt{q}}+z^{\sqrt{q}+1}=0.$
+Put $a_{21}a_{12}^{-{\sqrt{q}}}=\omega.$
+To confirm the last assertion is not difficult, actually $(1,0,0)$ and $(0,1,0)$ are inflexions. ∎
+By Corollary 3.7, we have done the proof of the first paragraph of Theorem 3.1.
+Now we investigate the curve (5), hereafter, this curve will be denoted by $C_{\omega}$.
+Lemma 3.8.
+The number of $\mathbb{F}_{q}$-points on $C_{\omega}\cap\{z=0\}$ is $\sqrt{q}+1$ if $\operatorname{Nm}\omega=1$, and $2$ if $\operatorname{Nm}\omega\neq 1$.
+Proof.
+For any $\omega$, $P=(1,0,0)$ and $Q=(0,1,0)$ are rational inflexions on $C_{\omega}$.
+Since $\{z=0\}\cap\{y=0\}=\{P\}$, it is enough to count the number of $\mathbb{F}_{q}$-solutions of (5) when $y=1$ and $z=0$.
+The equation in question is $x^{\sqrt{q}}+\omega x=0.$ Except for $Q$, the other points of $C_{\omega}\cap\{z=0\}\cap\{y=1\}$ correspond to the roots of $x^{\sqrt{q}-1}+\omega=0.$ By Corollary 3.3, the number of $\mathbb{F}_{q}$-solutions of this equation is
+$\left\{\begin{array}[]{cl}\sqrt{q}-1&\text{if $\operatorname{Nm}\omega\,(=\operatorname{Nm}(-\omega))=1$}\\ 0&\text{ else.}\end{array}\right.$
+Adding $2$ coming from $\{P,Q\}$ in each cases, we have the desired numbers. ∎
+Next we compute the number of rational points of $C_{\omega}$ outside of $\{z=0\}$.
+Lemma 3.9.
+The number of $\mathbb{F}_{q}$-solutions of affine equation
+$x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1=0$ (15)
+is
+$\left\{\begin{array}[]{cl}\sqrt{q}(q-1)&\text{ if $\omega=1$}\\ 0&\text{ if $\operatorname{Nm}\omega=1$ and $\omega\neq 1$}\\ q-1&\text{ if $\operatorname{Nm}\omega\neq 1$}.\end{array}\right.$
+Proof.
+If the affine equation (15) has an $\mathbb{F}_{q}$-solution $(x_{0},y_{0})$, then $x_{0}y_{0}^{\sqrt{q}}$ is a root of the equation in $T$
+$T^{\sqrt{q}}+\omega T+1=0.$ (16)
+Conversely if $t_{0}$ is a root of (16), then $(\frac{t_{0}}{y_{0}^{\sqrt{q}}},y_{0})$ is a solution of (15) for any $y_{0}\in\mathbb{F}_{q}^{\ast}$. Therefore the number of solutions of (15) is
+[…]
+Suppose that $\operatorname{Nm}\omega=1$.
+Then, by Proposition 3.4 (i), (16) has an $\mathbb{F}_{q}$-solution if and only if $\omega=1$.
+In this case, it has exactly $\sqrt{q}$ rational solutions, and totally (15) has $\sqrt{q}(q-1)$ rational solutions.
+Suppose that $\operatorname{Nm}\omega\neq 1$. Then, from the second part of the same proposition, (16) has a unique $\mathbb{F}_{q}$-solution, and then (15) has $(q-1)$ rational solutions. ∎
+The following three lemmas will be used for showing the part (b) of Theorem 3.1.
+The element of $PGL(3,\mathbb{F}_{q})$ which comes from the diagonal matrix with entries $a,b,c$ from the upper left to the lower right will be denoted by $\operatorname{diag}[a,b.c]$.
+Lemma 3.10.
+For $\eta\in\mathbb{F}_{q}^{\ast}$, we have
+$\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\eta\end{pmatrix}\sim\operatorname{diag}[1,1,\eta].$
+Proof.
+Choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$.
+Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then by straightforward computation, we know that $T_{0}^{\ast}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\eta].$
+Lemma 3.11.
+Let $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$.
+Then the number of rational points of $C_{\operatorname{diag}[1,1,\eta]}$ is $\sqrt{q}+1$, and these rational points are on the line $\{z=0\}$.
+Proof.
+The curve $C_{\operatorname{diag}[1,1,\eta]}$ is defined by
+$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0.$ (17)
 If $(x_{0},y_{0},z_{0})\in(\mathbb{F}_{q})^{3}$ is a solution of (17), then it is a solution of the simultaneous equations:
 […]
-because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}\in\mathbb{F}_{\sqrt{q}}$ and $\{1,\omega\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
-$D_{\omega}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}.$
-Hence $N_{q}(C_{A})=\sqrt{q}+1$, and $m_{C_{A}}=1$.
-Example 3.5
-(Type 2) $A\sim\operatorname{diag}[1,\omega,\omega^{\prime}]$ for some $\omega,\omega^{\prime}\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$ with $\omega(\omega^{\prime})^{-1}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
-Hence $C_{A}$ is projectively equivalent the curve $D_{\omega,\omega^{\prime}}$ defined by
-$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+\omega^{\prime}z^{\sqrt{q}+1}=0.$ (10)
-Since any two elements of $\{1,\omega,\omega^{\prime}\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, there is a unique relation up to $\mathbb{F}_{\sqrt{q}}^{\ast}$-multiple among those three elements, say
-$\alpha+\beta\omega+\gamma\omega^{\prime}=0.$
-Since none of three coefficients $\alpha,\beta,\gamma$ are 00, we may assume that $\alpha=1$, and the first coordinate of any solution of (10) is also $1$.
-Then $D_{\omega,\omega^{\prime}}(\mathbb{F}_{q})=\{(1,\zeta,\eta)|\operatorname{Nm}\zeta=\beta,\operatorname{Nm}\eta=\gamma\}$.
-Hence $N_{q}(C_{A})=(\sqrt{q}+1)^{2}$, and hence $m_{C}=\sqrt{q}+2$.
-4 Example II - non-diagonal case
-The following examples, including the diagonal case explained above, are inspired by Hirschfeld’s classification of cubic plane curves over any finite fields [2, Section 11].
-Actually, the family of supersingular plane cubic curves over $\mathbb{F}_{4}$ coincides with the family of relatives of the Hermitian curve over $\mathbb{F}_{4}$, because any curves in each family is characterized by disappearance of the term $xyz$ in its equation (see, [1, IV, Prop.
-4.21]).
-Example 4.1
-Let $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, and
-$A=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}.$
-Then $C_{A}$ is defined by
-$x^{\sqrt{q}}y+xy^{\sqrt{q}}+\omega z^{\sqrt{q}+1}=0.$
-The curve $C_{A}$ is projectively equivalent over $\mathbb{F}_{q}$ to a diagonal curve of type 1.
-In fact, choose $u\in\mathbb{F}_{q}$ so that $\operatorname{Tr}u=1$ and $\alpha\in\mathbb{F}_{q}^{\ast}$ so that $\operatorname{Nm}\alpha=-1$.
-Let $T_{0}=\begin{pmatrix}u&\alpha u^{\sqrt{q}}&0\\ 1&-\alpha&0\\ 0&0&1\end{pmatrix}.$ Then it is easy to see that
-${}^{t}\!T_{0}^{(\sqrt{q})}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&\omega\end{pmatrix}T_{0}=\operatorname{diag}[1,1,\omega].$
-Example 4.2
-Let us consider the matrix $A=\begin{pmatrix}1&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$. So the curve $C_{A}$ is defined by
-$x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (11)
-Proposition 4.3
-For the curve (11), $C_{A}(\mathbb{F}_{q})=\{(0,0,1)\}$, and hence $m_{C_{A}}=0$.
+because $x_{0}^{\sqrt{q}+1},y_{0}^{\sqrt{q}+1},z_{0}^{\sqrt{q}+1}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{1,\eta\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$. Therefore
+$C_{\operatorname{diag}[1,1,\eta]}(\mathbb{F}_{q})=\{(\zeta,1,0)|\operatorname{Nm}\zeta=-1\}\subset\{z=0\}.$
+This completes the proof. ∎
+Lemma 3.12.
+Let $\eta,\eta^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$. Then there is an element $T\in PGL(3,\mathbb{F}_{q})$ such that
+$T^{\ast}\operatorname{diag}[1,1,\eta]T=\operatorname{diag}[1,1,\eta^{\prime}]$
+if and only if $\eta^{\prime}\eta^{-1}\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
 Proof.
-For any $(x_{0},y_{0},z_{0})\in C_{A}(\mathbb{F}_{q})$, three elements $x_{0}^{\sqrt{q}+1}$, $y_{0}^{\sqrt{q}+1}$, and $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})$ are in $\mathbb{F}_{\sqrt{q}}$.
-Since $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, we have
-$x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=0\textrm{\ and \ }y_{0}^{\sqrt{q}+1}=0.$
-Hence $x_{0}=y_{0}=0$. ∎
-Example 4.4
-Next we consider the matrix $A=\begin{pmatrix}\omega&0&0\\ 0&\omega&1\\ 0&1&0\end{pmatrix},$ where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$.
-The corresponding curve $C_{A}$ is defined by
-$\omega x^{\sqrt{q}+1}+\omega y^{\sqrt{q}+1}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$ (12)
-Proposition 4.5
-For the curve (12),
-$\begin{split}C_{A}(\mathbb{F}_{q})=\{P_{\zeta}=(\zeta,1,0)|&\zeta^{\sqrt{q}+1}=-1\}\cup\{R=(0,0,1)\}\\ &\cup\{Q_{\xi\rho}=(\xi,\rho,1)|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.\end{split}$
-In particular, $N_{q}(C_{A})=q+\sqrt{q}+1$ and $m_{C_{A}}=\sqrt{q}+1$.
-Proof.
-On the line $\{z=0\}$ [^1] During computation of examples, frequently use this type of abbreviation., the equation is $\omega(x^{\sqrt{q}+1}+y^{\sqrt{q}+1})=0.$ Hence
-$C_{A}(\mathbf{F}_{q})\cap\{z=0\}=\{P_{\zeta}|\zeta^{\sqrt{q}+1}=-1.\}$
-In affine space $\mathbb{P}^{2}\setminus\{z=0\}$, we may assume that the $z$-coordinate is always $1$. If $(x_{0},y_{0},1)$ is an $\mathbb{F}_{q}$-point of $C_{A}$, it satisfies
-$\omega(x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1})+y_{0}^{\sqrt{q}}+y_{0}=0,$ (13)
-and vice versa.
-Since $x_{0}^{\sqrt{q}+1}+y_{0}^{\sqrt{q}+1}$ and $y_{0}^{\sqrt{q}}+y_{0}$ are elements of $\mathbb{F}_{\sqrt{q}}$, and $\{\omega,1\}$ are linearly independent over $\mathbb{F}_{\sqrt{q}}$, (13) is equivalent to
+Let $\lambda=\eta^{\prime}\eta^{-1}$, and suppose that $\lambda\in\mathbb{F}_{\sqrt{q}}^{\ast}$.
+Since $\operatorname{Nm}$ is surjective, there is an element $\alpha\in\mathbb{F}_{q}^{\ast}$ such that $\alpha^{\sqrt{q}+1}=\lambda$.
+So
+$\operatorname{diag}[1,1,\alpha]^{\ast}\operatorname{diag}[1,1,\eta]\operatorname{diag}[1,1,\alpha]=\operatorname{diag}[1,1,\eta^{\prime}].$
+Conversely, let $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
+$T^{\ast}\operatorname{diag}[1,1,\eta]T=\rho\operatorname{diag}[1,1,\eta^{\prime}],$
+where we are regarding matrices as in $GL(3,\mathbb{F}_{q})$ and $\rho\in\mathbb{F}_{q}^{\ast}$.
+For both curves $C_{\operatorname{diag}[1,1,\eta]}$ and $C_{\operatorname{diag}[1,1,\eta^{\prime}]}$, their sets of $\mathbb{F}_{q}$-points consist of $\sqrt{q}+1$ points on the line $\{z=0\}$ by Lemma 3.11.
+Hence the line $\{z=0\}$ is stable by the linear transformation of $\mathbb{P}^{2}$ corresponding $T$, that is
+$\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ a_{31}&a_{32}&a_{33}\end{pmatrix}\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}=\begin{pmatrix}\ast\\ \ast\\ 0\end{pmatrix}.$
+Hence $a_{31}=a_{32}=0$, and
+$\begin{pmatrix}a_{11}^{\sqrt{q}}&a_{21}^{\sqrt{q}}&0\\ a_{12}^{\sqrt{q}}&a_{22}^{\sqrt{q}}&0\\ a_{13}^{\sqrt{q}}&a_{23}^{\sqrt{q}}&a_{33}^{\sqrt{q}}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta\end{pmatrix}\begin{pmatrix}a_{11}&a_{12}&a_{13}\\ a_{21}&a_{22}&a_{23}\\ 0&0&a_{33}\end{pmatrix}=\rho\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&\eta^{\prime}\end{pmatrix}.$
+Comparing $(1,1)$-component of the both sides, we have $a_{11}^{\sqrt{q}+1}+a_{21}^{\sqrt{q}+1}=\rho$, which guarantees that $\rho$ is an element of $\mathbb{F}_{\sqrt{q}}^{\ast}$.
+Taking determinant of the both sides, we have
+$\det T^{(\sqrt{q})}\cdot\eta\cdot\det T=\rho^{3}\eta^{\prime}.$
+Hence $\eta^{\prime}\eta^{-1}=(\det T)^{\sqrt{q}+1}\rho^{-3}\in\mathbb{F}_{\sqrt{q}}^{\ast}$
+Proof of Theorem 3.1.
+(a) Suppose that $\omega=1$.
+Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}^{3}+1$ rational points.
+Hence it is Hermitian if $q\neq 4$ [1], and all rational points on a Hermitian curve are inflexions.
+When $q=4$, a non-Hermitian curve with $9$ rational points can’t have inflexions [3, Remark 3.10].
+(b) Suppose that $\operatorname{Nm}\omega=1$ and $\omega\neq 1$.
+Then by Lemmas 3.8, 3.9, the curve $C_{\omega}$ has $\sqrt{q}+1$ rational points, and those points lie on the line $\{z=0\}$.
+Let $R$ be one of those points.
+If $R$ is not an inflexion, then $C_{\omega}.T_{R}(C_{\omega})=\sqrt{q}R+R^{\prime}$ with $R^{\prime}\neq R$.
+Since $R^{\prime}$ is $\mathbb{F}_{q}$-point too, $R^{\prime}$ lies on the line $\{z=0\}$.
+Hence $T_{R}(C_{\omega})$ coincides with the line $\{z=0\}$, which is impossible.
+Therefore all the $\sqrt{q}+1$ rational points are inflexions.
+Next we show that $C_{\omega}$ with $\operatorname{Nm}\omega=1$ and $\omega\neq 1$ is projectively equivalent to
+$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\eta z^{\sqrt{q}+1}=0$ (18)
+for a certain $\eta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}.$ Since $\operatorname{Nm}\omega=1$, there is an element $\beta\in\mathbb{F}_{q}^{\ast}\setminus\mathbb{F}_{\sqrt{q}}^{\ast}$ such that $\beta^{1-\sqrt{q}}=\omega$.
+Using this $\beta$, (5) can be rewritten as
+$(\beta x)^{\sqrt{q}}y+(\beta x)y^{\sqrt{q}}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$ (19)
+Replacing $\beta x$ with the new $x$ and applying Lemma 3.10, we get the equation
+$x^{\sqrt{q}+1}+y^{\sqrt{q}+1}+\beta^{\sqrt{q}}z^{\sqrt{q}+1}=0.$
+Replace $\beta^{\sqrt{q}}$ by $\eta$, then we have (18). Conversely, tracing the reverse process, we know that the curve defined by (18) is projectively equivalent to $C_{\eta^{\sqrt{q}-1}}$.
+The last statement in (b) is just Lemma 3.12.
+(c) Suppose that $\operatorname{Nm}\omega\neq 1$.
+From Lemmas 3.8, 3.9, $C_{\omega}(\mathbb{F}_{q})$ consists of $q+1$ rational points.
+Note that $P=(1,0,0)$ and $Q=(0,1,0)$ are inflexions, and they are only $\mathbb{F}_{q}$-points lie on the line $\{z=0\}$ by Lemma 3.5.
+We will show that any point $R\in C_{\omega}(\mathbb{F}_{q})\setminus\{P,Q\}$ is not an inflexion.
+Since $R$ does not lie on the line $\{z=0\}$, we may assume that $R=(x_{0},y_{0},1)$, where $(x_{0},y_{0})$ satisfies the affine equation (15), in particular $x_{0}\neq 0$.
+We want compute the divisor $C_{\omega}.T_{R}(C_{\omega})$.
+Let $f(x,y)=x^{\sqrt{q}}y+\omega xy^{\sqrt{q}}+1$.
+Since $f_{x}=\omega y^{\sqrt{q}}$ and $f_{y}=x^{\sqrt{q}}$, the tangent line at $R$ is given by
+$\omega y_{0}^{\sqrt{q}}x+x_{0}^{\sqrt{q}}y+1=0,$
+or, alternatively
+$y=-\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}x-(\frac{1}{x_{0}})^{\sqrt{q}}$ (20)
+because $x_{0}\neq 0$. Take $\sqrt{q}$-th power of (20), then
+$y^{\sqrt{q}}=-\omega^{\sqrt{q}}(\frac{y_{0}}{x_{0}})x^{\sqrt{q}}-(\frac{1}{x_{0}}).$ (21)
+Substitute (20) and (21) into (15), then
+$-\left\{\left(\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}\right)x^{\sqrt{q}+1}+(\frac{1}{x_{0}})^{\sqrt{q}}x^{\sqrt{q}}+\frac{\omega}{x_{0}}x-1\right\}=0.$ (22)
+On the other hand, since
+$0=\omega\left(\frac{f(x_{0},y_{0})}{x_{0}^{\sqrt{q}+1}}\right)^{\sqrt{q}}=\omega(\frac{y_{0}}{x_{0}})^{\sqrt{q}}+\omega^{\sqrt{q}+1}\frac{y_{0}}{x_{0}}+\frac{\omega}{x_{0}^{\sqrt{q}+1}},$
+the coefficient of $x^{\sqrt{q}+1}$ in (22) is $\frac{\omega}{x_{0}^{\sqrt{q}+1}}$. Hence (22) can be written as
+$\frac{\omega}{x_{0}^{\sqrt{q}+1}}x^{\sqrt{q}+1}-\frac{1}{x_{0}^{\sqrt{q}}}x^{\sqrt{q}}-\frac{\omega}{x_{0}}x+1$ (23)
+$=$ $\frac{\omega}{x_{0}^{\sqrt{q}+1}}\left(x-x_{0}\right)^{\sqrt{q}}\left(x-\frac{x_{0}}{\omega}\right)=0.$ (24)
+Since $\omega\neq 1$, $x_{0}\neq\frac{x_{0}}{\omega}$, which means that $R$ is not an inflexion. Hence only $P$ and $Q$ are inflexions of $C_{\omega}$.
+Lastly we classify the curves $C_{\omega}$ with $\omega\neq 1$ up to projective equivalence. For $\omega,\omega^{\prime}\in\mathbb{F}_{q}^{\ast}\setminus\operatorname{Ker}(\operatorname{Nm})$, let
+$A_{\omega}=\begin{pmatrix}0&1&0\\ \omega&0&0\\ 0&0&1\end{pmatrix}\text{ and }A_{\omega^{\prime}}=\begin{pmatrix}0&1&0\\ \omega^{\prime}&0&0\\ 0&0&1\end{pmatrix}.$
+Then two curves $C_{\omega}$ and $C_{\omega^{\prime}}$ are projectively equivalent to each other if and only if there is a matrix $T=(a_{ij})\in GL(3,\mathbb{F}_{q})$ such that
+$T^{\ast}A_{\omega}T=\rho A_{\omega^{\prime}}$ (25)
+for some $\rho\in\mathbb{F}_{q}^{\ast}$. Since inflexions of both curves are $\{P,Q\}$ that are lie on the line $\{z=0\}$, and
+$T_{P}(C_{\omega})\cap T_{Q}(C_{\omega})=T_{P}(C_{\omega^{\prime}})\cap T_{Q}(C_{\omega^{\prime}})=\{(0,0,1)\},$
+the line $\{z=0\}$ and the point $(0,0,1)$ are stable by the projective transformation $T$.
+Hence we may choose $T$ as $\begin{pmatrix}a_{11}&a_{12}&0\\ a_{21}&a_{22}&0\\ 0&0&1\end{pmatrix}.$ Then, comparing $(3,3)$-entries of both sides of (25), we know $\rho=1$.
+Now we write down (25) componentwise:
+$\omega a_{21}^{\sqrt{q}}a_{11}+a_{11}^{\sqrt{q}}a_{21}$ $=0$ (26)
+$\omega a_{22}^{\sqrt{q}}a_{12}+a_{12}^{\sqrt{q}}a_{22}$ $=0$ (27)
+$\omega a_{21}^{\sqrt{q}}a_{12}+a_{11}^{\sqrt{q}}a_{22}$ $=1$ (28)
+$\omega a_{22}^{\sqrt{q}}a_{11}+a_{12}^{\sqrt{q}}a_{21}$ $=\omega^{\prime},$ (29)
+where (26)–(29) come from comparison of $(1,1)$,(2,2)22(2,2)( 2 , 2 ), $(1,2)$ and $(2,1)$ components respectively. Take $\sqrt{q}$th power of (26)
+$\omega^{\sqrt{q}}a_{21}a_{11}^{\sqrt{q}}+a_{11}a_{21}^{\sqrt{q}}=0,$ (30)
 […]
-Hence
-$C_{A}(\mathbb{F}_{q})\cap\{z\neq 0\}=\{R\}\cup\{Q_{\xi\rho}|\rho^{\sqrt{q}-1}=-1,\ \xi^{\sqrt{q}+1}=\rho^{2}\}.$
-It is easy to show that |$|\{P_{\xi}\}|=\sqrt{q}+1$, and |$|\{R\}\cup\{Q_{\xi\rho}\}|=1+(\sqrt{q}-1)(\sqrt{q}+1).$
-Example 4.6
-Let $A=\begin{pmatrix}1&0&0\\ 1&0&1\\ 0&1&0\end{pmatrix},$ Hence the curve $C_{A}$ is defined by
-$x^{\sqrt{q}+1}+xy^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=0.$
-Proposition 4.7
-For the curve $C_{A}$ in Example 4.6, we have
-$N_{q}(C_{A})=q+1,\text{\ and so\ }m_{C_{A}}=\sqrt{q}.$
-Proof.
-First consider the case $x=0$. Then the equation becomes $y^{\sqrt{q}}z+yz^{\sqrt{q}}=0$, which has obviously $\sqrt{q}+1$ homogeneous solutions that are $\mathbb{F}_{q}$points.
-Next suppose that $x=1$. Then the equation becomes
-$01+y^{\sqrt{q}}+y^{\sqrt{q}}z+yz^{\sqrt{q}}=$ 0 . (14)
-Let $(y_{0},z_{0})$ be any $\mathbb{F}_{q}$-solution of (14).
-Since $y_{0}^{\sqrt{q}}z_{0}+y_{0}z_{0}^{\sqrt{q}}=\operatorname{Tr}(y_{0}^{\sqrt{q}}z_{0})\in\mathbb{F}_{\sqrt{q}}$, the remaining term $1+y_{0}^{\sqrt{q}}$ is also an element of $\mathbb{F}_{\sqrt{q}}$, and so is $y_{0}^{\sqrt{q}}$.
-Hence $y_{0}=y_{0}^{q}=(y_{0}^{\sqrt{q}})^{\sqrt{q}}\in\mathbb{F}_{\sqrt{q}}$.
-Therefore the point $(y_{0},z_{0})$ satisfies
-$01+y_{0}+(z_{0}+z_{0}^{\sqrt{q}})y_{0}=$ 0 (15)
-from (14).
-Put $z_{0}+z_{0}^{\sqrt{q}}=t_{0}$, which is an elements of $\mathbb{F}_{\sqrt{q}}$.
-If $t_{0}\neq-1$, $y_{0}=\frac{-1}{t_{0}+1}$, which is uniquely determined by $t_{0}$.
-For each $t_{0}\in\mathbb{F}_{\sqrt{q}}\setminus\{-1\}$, we can find $\sqrt{q}$ solutions of $z+z^{\sqrt{q}}=t_{0}$ in $\mathbb{F}_{q}$.
-If $t_{0}=-1$, There are no solutions of (15) on $y_{0}$.
-Hence $|C_{A}(\mathbb{F}_{q})\cap\{x\neq 0\}|=\sqrt{q}(\sqrt{q}-1).$ Therefore the number of $\mathbb{F}_{q}$-points of $C_{A}$ is totally
-$(\sqrt{q}+1)+\sqrt{q}(\sqrt{q}-1)=q+1.$
-This completes the proof. ∎
-5 Example III - non-diagonal case (2)
-In this section, we handle the curve
-$x^{\sqrt{q}}z+xy^{\sqrt{q}}+\omega yz^{\sqrt{q}}=0,$ (16)
-where $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$.
-The number of $\mathbb{F}_{q}$-points of curves in each examples explained in previous two sections does not depend on the choice of $\omega\in\mathbb{F}_{q}\setminus\mathbb{F}_{\sqrt{q}}$, however that of the curve (16) depends on the choice of $\omega$.
-So we denote the corresponding matrix $\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&\omega&0\end{pmatrix}$ by $A_{\omega}$.
-Lemma 5.1
-Let $C_{A_{\omega}}$ be a curve defined by (16).
-(i)   $C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z=0\}=\{(1,0,0),(0,1,0)\}.$
-${\rm(ii)}\begin{split}(C_{A_{\omega}}(\mathbb{F}_{q})\cap&\{z\neq 0\})\cap\{y^{\sqrt{q}+1}\neq 1\}\\ &=\left\{\left(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1},y_{0},1\right)\mid y_{0}\in\mathbb{F}_{q},\ y_{0}^{\sqrt{q}+1}-1\neq 0\right\}.\end{split}$
-In particular,
-$N_{q}(C_{A_{\omega}})\geq q-\sqrt{q}+1.$
-Proof.
-The assertion (i) is obvious. Let $(x_{0},y_{0},1)\in C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\}$. Then
-$x_{0}^{\sqrt{q}}+x_{0}y_{0}^{\sqrt{q}}+\omega y_{0}=0.$ (17)
-Taking $\sqrt{q}$-th power of this equation, we have
-$x_{0}+x_{0}^{\sqrt{q}}y_{0}+\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}=0.$ (18)
-[…]
-$x_{0}(y_{0}^{\sqrt{q}+1}-1)-(\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2})=0.$
-Hence, when $y_{0}^{\sqrt{q}+1}-1\neq 0$, $x_{0}=\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1}$. Conversely, by straightforward computation, we get
-$(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})^{\sqrt{q}}+(\frac{\omega^{\sqrt{q}}y_{0}^{\sqrt{q}}-\omega y_{0}^{2}}{y_{0}^{\sqrt{q}+1}-1})y_{0}^{\sqrt{q}}+\omega y_{0}=0.$
-Putting (i), (ii) together, we have $q-(\sqrt{q}+1)+2$ points over $\mathbb{F}_{q}$. ∎
-For computing the number of $\mathbb{F}_{q}$-points of $(C_{A_{\omega}}(\mathbb{F}_{q})\cap\{z\neq 0\})\cap\{y^{\sqrt{q}+1}=0\},$ we fix a primitive element $\omega_{0}\in\mathbb{F}_{q}^{\ast}$.
-Note that $\omega_{0}^{s}\in\mathbb{F}_{\sqrt{q}}^{\ast}$ if and only if $s\equiv 0\mod\sqrt{q}+1$.
-Lemma 5.2
-(1) Suppose that $q$ is odd. Then
-(i) $-1=\omega_{0}^{\frac{q-1}{2}},$ and
-(ii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{\frac{\sqrt{q}+1}{2}+s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}.$
-(2) Suppose that $q$ is even. Then
-(iii) $\ker\operatorname{Tr}=\{0\}\cup\{\omega_{0}^{s(\sqrt{q}+1)}\mid s=0,1,\dots,\sqrt{q}-2\}$.
-Proof.
-(i) Since $\omega_{0}$ is of order $q-1$ in the multiplicative group $\mathbb{F}_{q}^{\ast}$ and $q-1$ is even, $\omega_{0}^{\frac{q-1}{2}}=-1$.
-(ii) For $b\in\mathbb{F}_{q}^{\ast}$, $\operatorname{Tr}b=0$ means $b^{\sqrt{q}-1}=-1$.
-Letting $b=\omega_{0}^{l}$, we have $\omega_{0}^{l(\sqrt{q}-1)}=\omega_{0}^{\frac{q-1}{2}}$.
-Hence $l(\sqrt{q}-1)\equiv\frac{\sqrt{q}+1}{2}(\sqrt{q}-1)\mod q-1$, which means that $l\equiv\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1$.
-(iii) Similar to the case (ii). ∎
-Theorem 5.3
-  1. (1)
-    If $(\sqrt{q}+1,3)=1$, then
-    $N_{q}(C_{A_{\omega}})=q+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}.$
-  2. (2)
-    When $\sqrt{q}+1$ is divisible by $3$,
-    1. (a)
-      if $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
-      $N_{q}(C_{A_{\omega}})=q+2\sqrt{q}+1\text{\ and so \ }m_{C_{A_{\omega}}}=\sqrt{q}+2,\text{\ and}$
-    2. (b)
-      if $\omega^{\frac{\sqrt{q}+1}{3}}\not\in\mathbb{F}_{\sqrt{q}}^{\ast}$, then
-      $N_{q}(C_{A_{\omega}})=q-\sqrt{q}+1\text{\ and \ }m_{C_{A_{\omega}}}=\sqrt{q}-1.$
-Proof.
-We have to count the number of $\mathbb{F}_{q}$-solutions of the simultaneous equations
-[…]
-By changing the sign $y\mapsto-y,$ this system is equivalent to
-[…]
-For a given $y_{0}\in\mathbb{F}_{q}^{\ast}$ with $y_{0}^{\sqrt{q}+1}=1$, we ask whether the equation on $x$
-$x^{\sqrt{q}}-y_{0}^{\sqrt{q}}x-\omega y_{0}=0$ (19)
-has an $\mathbb{F}_{q}$-solution or not.
-Since $\operatorname{Nm}y_{0}^{\sqrt{q}}=\operatorname{Nm}y_{0}=1$, there is an element $c\in\mathbb{F}_{q}^{\ast}$ such that $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$ by Lemma 3.1.
-Therefore the problem above is transformed whether the equation on $cx$
-$(cx)^{\sqrt{q}}-(cx)-\omega c^{\sqrt{q}}y_{0}=0$ (20)
-has an $\mathbb{F}_{q}$-solution or not.
-From Corollary 3.2, (20) has an $\mathbb{F}_{q}$-solution if and only if $\operatorname{Tr}(\omega c^{\sqrt{q}}y_{0})=0$, and in this case all solutions of (20) are $\mathbb{F}_{q}$-solutions.
-Let $i,j,k$ be integers with $11\leq i,j,k\leq q-$ 1 satisfying
-$\omega=\omega_{0}^{k},\ c=\omega_{0}^{j},\text{\ and \ }y_{0}=\omega_{0}^{i}.$
-Since $y_{0}^{\sqrt{q}+1}=1$, there is a unique integer $l$ with $11\leq l\leq\sqrt{q}+$ 1 so that $i\equiv l(\sqrt{q}-1)\mod q-1$. Since $c^{1-\sqrt{q}}=y_{0}^{\sqrt{q}}$,
-$j(1-\sqrt{q})\equiv l(\sqrt{q}-1)\sqrt{q}\mod q-1.$
-Hence
-$j\equiv-l\sqrt{q}\mod\sqrt{q}+1.$ (21)
-Since $\omega c^{\sqrt{q}}y_{0}=\omega_{0}^{k+j\sqrt{q}+l(\sqrt{q}-1)},$ $\operatorname{Tr}\omega c^{\sqrt{q}}y_{0}=0$ if and only if
-$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}\end{cases}$ (22)
-by Lemma 5.2. On the other hand,
-$k+j\sqrt{q}+l(\sqrt{q}-1)\equiv k-l(q-\sqrt{q}+1)\mod\sqrt{q}+1$ (23)
-by (21). Together with $\sqrt{q}\equiv-1\mod\sqrt{q}+1$ and $q\equiv-\sqrt{q}\equiv 1\mod\sqrt{q}+1,$ (22) and (23) imply
-$k-3l\equiv\begin{cases}\frac{\sqrt{q}+1}{2}\mod\sqrt{q}+1&\text{ if $q$ is odd}\\ 0\mod\sqrt{q}+1&\text{ if $q$ is even}.\end{cases}$ (24)
-Therefore the number of $y_{0}$’s for each of which (19) has $\mathbb{F}_{q}$-solutions, in fact $\sqrt{q}$-solutions, coincides with the number of $l$’s with $11\leq l\leq\sqrt{q}+$ 1 each of which satisfies (24) for the fixed $k$.
-Let us consider the map
-$\begin{array}[]{rccc}\varphi:&\mathbb{Z}/(\sqrt{q}+1)&\to&\mathbb{Z}/(\sqrt{q}+1)\\ &x&\mapsto&3x\end{array}.$
-If $\sqrt{q}+1$ is not divisible by $3$, $\varphi$ is bijective. Hence there exists a unique $l$ satisfying (24) for any fixed $k$. If $3$ is a factor of $\sqrt{q}+1$, then
-$|\varphi^{-1}(m)|=\begin{cases}3&\text{if $m\in\operatorname{Im}\varphi$}\\ 0&\text{else}.\end{cases}$
-Hence if $k\equiv 0\mod 3$, there are three $l$’s satisfying (24), and otherwise no $l$’s.
-The condition $k\equiv 0\mod 3$ is equivalent to the condition $\omega^{\frac{\sqrt{q}+1}{3}}\in\mathbb{F}_{\sqrt{q}}^{\ast}.$ This completes the proof.
-∎
-Finally we pose a question: when $q\geq 16$, does a Hermitian-relative curve $C$ with $22\leq m_{C}\leq\sqrt{q}-$ 2 exist?
+$(\omega^{\sqrt{q}+1}-1)a_{11}^{\sqrt{q}}a_{21}=0.$
+Since $\operatorname{Nm}\omega\neq 1$, we conclude that $a_{11}=0$ or $a_{21}=0$. Similarly, using (27), we have $a_{22}=0$ or $a_{12}=0$. Since $\det T\neq 0$, only possibilities of $T$ are
+$T=\begin{pmatrix}a_{11}&0&0\\ 0&a_{22}&0\\ 0&0&1\end{pmatrix}\text{or}\begin{pmatrix}0&a_{12}&0\\ a_{21}&0&0\\ 0&0&1\end{pmatrix}.$
+If the first case occurs, then $a_{11}^{\sqrt{q}}a_{22}=1$ by (28) and $\omega a_{22}^{\sqrt{q}}a_{11}=\omega^{\prime}$ by (29), which implies that $\omega=\omega^{\prime}$.
+If the second case occurs, then
+$\omega a_{21}^{\sqrt{q}}a_{12}=1\text{ and }a_{12}^{\sqrt{q}}a_{21}=\omega^{\prime},$
+which implies that $\omega\omega^{\prime\sqrt{q}}=1$. ∎
 Acknowledgment. The second author was partially supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (2022R1A2C1012291).